OpenStudy (unklerhaukus):

Solve the DE\[yy''+(y')^2=0\]

OpenStudy (unklerhaukus):

the equation is x-absent

OpenStudy (unklerhaukus):

\[\text{let }\frac{\text d y}{\text dx}=p\]\[\frac{d^2y}{dx^2}=p\frac{dp}{\text d y}\]

OpenStudy (unklerhaukus):

\[yp\frac{\text d p}{\text dy}-p^2=0\]

OpenStudy (unklerhaukus):

\[\frac{\text dp}{\text dy}-\frac1yp=0\]

OpenStudy (unklerhaukus):

\[I.F. (y)=e^{\int \frac 1y \text dy}=e^{\ln y}=y\]

OpenStudy (unklerhaukus):

\[\frac{\text d}{\text dy}\left(py\right)=0\]\[p=0\]

OpenStudy (unklerhaukus):

???

OpenStudy (anonymous):

is the ans y=ae^(bx)

OpenStudy (unklerhaukus):

the back of my book neglect this particular question for some reason, how did you arrive at\[y=ae^{bx}\]?

OpenStudy (anonymous):

dividing both sides by y^2 we have y'=by and then move on further .... by the way if u don't mind name/author of the book?

OpenStudy (unklerhaukus):

that is the chapter of the text book i am working on and the solutions chapter

OpenStudy (unklerhaukus):

and this question i am working on at the momnet is from te problem set 3A 1.g)

OpenStudy (anonymous):

thanx...

OpenStudy (anonymous):

ur questions are different -------yy′′−(y′)2=0 may be just check the sign whether its - or +

OpenStudy (unklerhaukus):

Its just a booklet the Uni produces { University of Wollongong } the subject is MATH202 Differential Equations II i could up load any chapter or the entire text if you want

OpenStudy (unklerhaukus):

your right

OpenStudy (anonymous):

yy′′+(y′)2=0 might be u have posted a different question ...

OpenStudy (unklerhaukus):

i just changed it

OpenStudy (unklerhaukus):

\[yy''+(y')^2=0\]

OpenStudy (unklerhaukus):

sorry about that

OpenStudy (anonymous):

the way u were doing was correct ... it reduces to yy'=c and then i hope u can further solve...

OpenStudy (anonymous):

sometimes error happens no need to worry

OpenStudy (unklerhaukus):

\[y\times 0=0?\]

OpenStudy (unklerhaukus):

yeah the second mistake i made way to leave out the negative sign in the integrating factor step

OpenStudy (anonymous):

integrating constant is missing...

OpenStudy (unklerhaukus):

\[-p=c\] \[?

OpenStudy (anonymous):

didn't get ...

OpenStudy (unklerhaukus):

I get so confused in that chapter under section x-absent page 1 it says ..."...then, \[p=\frac{\text dy}{\text dx}=\frac{\text d^2y}{\text d x^2}\]" is this right?

OpenStudy (experimentx):

\[ (yy')' = 0 \] \[ yy' = k_1\] \[ \frac{y^2}{2} = k_1x + k_2\]

OpenStudy (anonymous):

i don't agree with the textbooks statement \[p=\frac{dy}{dx}=\frac{d^2y}{dx^2}\] since: \[\frac{d^2y}{dx^2}=\frac{d}{dx}(\frac{dy}{dx})=\frac{dp}{dx}=\frac{dp}{dy}\frac{dy}{dx}=\frac{dp}{dy}y^'=p \frac{dp}{dy}\] Also, the equation is seperable where you do IF.

OpenStudy (unklerhaukus):

i thought it was a typo too but i wasent sure

OpenStudy (anonymous):

looks like a pretty decent online text though

OpenStudy (unklerhaukus):

it is from the booklet i have it in hard copy too

OpenStudy (unklerhaukus):

two versions actually i failed this subject

OpenStudy (anonymous):

sorry i left in between that as i had some work.... i hope ur answer is solved